### v1

# uncompress

[…]
uncompress
Statistical laws and linguistics inform meaning
in naturalistic and fictional conversation
Ashley M. A. Fehr Computational Story Lab, Vermont Advanced Computing Center, University of Vermont, Burlington, VT 05405, US Calla G. Beauregard Computational Story Lab, Vermont Advanced Computing Center, University of Vermont, Burlington, VT 05405, US Vermont Complex Systems Institute, MassMutual Center of Excellence for Complex Systems and Data Science, University of Vermont, Burlington, VT 05405, US Julia Witte Zimmerman Computational Story Lab, Vermont Advanced Computing Center, University of Vermont, Burlington, VT 05405, US Vermont Complex Systems Institute, MassMutual Center of Excellence for Complex Systems and Data Science, University of Vermont, Burlington, VT 05405, US Computational Ethics Lab, University of Vermont, Burlington, VT 05405, US Katie Ekström Glass Brain Lab, Department of Biomedical Engineering; Vermont Conversation Lab, Larner College of Medicine, University of Vermont, Burlington, VT 05405, US Pablo Rosillo-Rodes Computational Story Lab, Vermont Advanced Computing Center, University of Vermont, Burlington, VT 05405, US Institute for Cross-Disciplinary Physics and Complex Systems IFISC (UIB-CSIC), Campus Universitat de les Illes Balears, E-07122 Palma de Mallorca, Spain Christopher M. Danforth Computational Story Lab, Vermont Advanced Computing Center, University of Vermont, Burlington, VT 05405, US Vermont Complex Systems Institute, MassMutual Center of Excellence for Complex Systems and Data Science, University of Vermont, Burlington, VT 05405, US Department of Mathematics & Statistics, University of Vermont, Burlington, VT 05405, US Peter Sheridan Dodds Computational Story Lab, Vermont Advanced Computing Center, University of Vermont, Burlington, VT 05405, US Vermont Complex Systems Institute, MassMutual Center of Excellence for Complex Systems and Data Science, University of Vermont, Burlington, VT 05405, US Santa Fe Institute, 1399 Hyde Park Rd, Santa Fe, NM 87501, US
(December 19, 2025)
[…]
1 Introduction
Language is social, conveying meaning and understanding between individuals and groups [clark1996using_lang]. We employ language in conversation to fulfill both inter- and intra-personal needs, supporting communication, collaboration, and social connection [baumeister1995needtobelong, benus2014social_entrainment]. In addition, how people engage in conversation can be linked to public health outcomes such as well-being [sun2020_wellbeing, wheatley2024], motivating the study of conversation at scale [holt2017soc_connection_priority]. However, conversation in general, and the nature of turn-taking specifically (i.e., who has the floor, when, and for how long), are understudied [cooney2024naturalturn]. Conversation participants mutually determine the trajectory of their exchange, weaving a complex dynamic story [soloman2023, carragher2024coconstructedcomm, clark1996using_lang] in which singular processes may be insufficient to describe outcomes  [paletz2023speaking_similarly]. Therefore, methods from the analysis of dynamical systems can help capture emergent properties that remain elusive at the individual scale.
Linguistic scaling measures offer a way to understand more about the shape of information and symbols that underlie the construction of meaning in conversation (Fig. 1). With linguistic scaling, we can explore how statistical patterns can quantify and ultimately illuminate social processes. Previous work on linguistic scaling (e.g., Heaps’ law and Zipf’s law) is dominated by long-form, monologic text such as the Google Books corpus [petersen2012langs_cool_as_they_expand, pechenick2017lexicalturbulencescaling] and literary works [chacoma_tags_2020, gerlach_vocabgrowth2013, bian2016scaling_like_lin, altmann2024statisticallaws, williams2015text_mixing_rank_freq, williams2016zipf_coherent_lang_production, williams2015zipfs_phrases], leaving open questions as to how prior observations extend to language in general and within conversational mediums in particular.
[…]
Figure 1: An analogy wherein information is to meaning as lemons are to lemonade. The highlighted word stem ‘scoop’ in the top and bottom conversation bubbles demonstrate that context changes the word’s meaning—the differing contexts being ice cream (top) compared to someone publishing your idea before you, i.e., being ‘scooped’ (bottom).
1.1 Conversation and medium
Interlocutors co-construct communication, incorporating context such as language choices, personal knowledge, and social roles and identities to negotiate shared meaning [carragher2024coconstructedcomm]. Conversation is co-constructed [carragher2024coconstructedcomm, ruhlemann2007ConversationContext] dialogue between two or more participants [yeomans2023guide_to_convo_research, yeomans2022convo_circumplex, brooks2025talk] where turn-taking [sacks1974turntaking_fundamental, stivers2009turntaking_universal] and its coordination [meyer2023timing_natural_convo, ruhlemann2007ConversationContext] is a fundamental feature. Some work defines conversation as a sub-genre of communication that is face-to-face, spoken, and synchronous, permitting investigation of acoustic and prosodic language features. Despite the persistence of turn-taking as fundamental, this perspective excludes live texting [azios_clinical_2023, azios2023textingconvos] and conversations in formal texts, such as novels. Another thread of conversation research focuses on clinical implications for individuals with medical conditions (e.g., conversation analysis in aphasia [azios_clinical_2023] and traumatic brain injury [azios2018convo_analysis_tbi]) or the qualities of patient-provider conversation [ross2020story_arcs_serious_illness]. This clinical work typically derives from face-to-face spoken conversation that is analyzed acoustically using audio or video signals or converted to written transcriptions.
Broadly, we construe conversation to include forms of dialogue or discourse that vary beyond the face-to-face, spoken modality. Thus, conversation substrates include spoken or naturalistic conversation, texting or direct message conversation, fictional or scripted conversation, and discussion forum conversation or microblogging. Conversation is also affected by the modality [boudin2024multimodal_feedback_in_convo] and the technology on which it transmits [williams1974tv_technology_culture, mcluhan1964medium], so our broad definition facilitates exploration across mediums in which conversational text plays different roles, and where the medium may facilitate those differences in conversation structure and function.
Conversation medium affects communication outcomes because of unique limitations and affordances of each substrate. Many mediums act as conversation substrates, for example, by defining the limitations on how much communication can occur (character limits or post frequency), which topics (moderation), how one can reply, and who (members) can participate in discussion. In defining these communication aspects, the medium shapes the structure of the conversation. In our research, we contrast naturalistic and scripted conversation specifically (summarized in Table 1). Naturalistic or spoken conversation has vocal and gestural features and is spontaneously structured. Scripted conversation includes dialogue that is edited or cognitively prepared, i.e., rehearsed, which can facilitate a clearer structure. Although scripted conversation’s written form might suggest it is more akin to other written texts, such as novels, we also see it as having some similarities with naturalistic conversation. Importantly, both substrates center the resolution of a speaker’s intended meaning and audience understanding [clark1996using_lang]. This shared context motivates us to compare underlying differences in the communication features of each medium and their potential impact on the construction of meaning. For these reasons, we expand on conversational data properties that differ by medium, and we consider how they shape conversation research.
Movie dialogue is delivered conversationally but created from non-participant perspectives. Instead, writer(s) must infer the part of the listener in order to structure the communication efficiently and with enough differentiating signal that the audience will follow along [linders2023zipf_spoken_dialog]. There is a semblance of turn co-construction, but all turns are ideated by the same third party at a time scale not captured by the final dialogue. Movie dialogue can also have additional layers of musical, cinematographic, and literary activity [clark1996using_lang] that scaffold the dialogue’s intended story development into each utterance [zago2020film_discourse, martin2010pos_bnc_movies]. Movie dialogues are situated as natural conversation, but scripts avoid features like overlapping speech and filler words, with some features more like those of literary works, such as having narrative purpose [zago2020film_discourse]. Together, these features of fictional conversations may reflect, at turns, both their unique role in storytelling and the qualities that people associate with natural conversation. Over narrative time, patterned use of communication features impacts the evolution and interpretation of narrative trajectory, offering insight into mental models of how conversation works. This contrast among features motivates our analysis of how conversational medium and attributes of the data itself (e.g., semantic content) shape both communication and the research seeking to describe it.
Naturalistic Scripted
Conception Voice (speaking, signing), spontaneously composed. Edited, speaker cognitive preparation, author does not equate to speaker.
Linguistic features Parallel speech, pauses, choice of word/tense/etc., audio. Derived from text first, written as intended.
Data
handling
Flattened information, diarizing, accuracy in turn parsing and transcribing, judgments about how to parse. Artificial lengths: turns, conversations, generally a turn-based structure that facilitates parsing.
Table 1: Conversation data contrast summary with example features of naturalistic and scripted modalities.
Considering a corpus’ features with respect to a naturalistic (versus non) comparison can also help determine the appropriate data handling. For example, spoken turns in naturalistic conversation might be longer, with more overlapping speech than in non-naturalistic conversation; assessing this aspect requires accurate turn segmentation and parallel speech processing. Lastly, the flattening of spoken conversation when transcribed (compared to text-based being instantiated as intended) can cause the loss of meaningful information, such as prosodic contours signaling pragmatic functions or overlapping speech that provides real-time feedback to the speaker on how their intended message is being received.
We compare naturalistic conversation and fictional conversation (dialogue exchanges in movie scripts) to tease apart how linguistic units differ, particularly through the lens of Heaps’ law and Zipf’s law, with attention to potential mechanisms for those differences (especially to the impact of spontaneity versus pre-planning).
1.2 Scaling in language
For text corpora, Heaps’ law classically describes a scaling relationship between the total number of words—tokens—and the number of unique words—types [peirce1906type_token_law, lu2010zipfandheaprelation]. For example, when reading through the first $t$ words of a book, the number of unique words $N(t)$ scales as $N(t)\sim t^{\beta}$ where $0\leq\beta\leq 1$. Types collectively comprise a text’s vocabulary $N(t)$. In general, a vocabulary may also contain non-words such as numbers and punctuation. More generally for complex systems, Heaps’ law is an instance of a type-entity relationship (e.g., tree species and individual trees).
Heaps’ law is understood to arise for growing systems which comprise many types that follow Zipf’s law [lu2010zipfandheaprelation, rosillo_rodes2025heap_paper]. For language, Zipf’s law is a power-law frequency-rank relationship expressed as $S_{r}\sim r^{-\alpha}$ where $r$ is the rank of any given word and $S_{r}$ is its count (and for general systems, $S$ indicates size). While the connection between Zipf’s law and Heaps’ law is not a simple scaling for finite systems, in the limit of large texts [rosillo_rodes2025heap_paper], $\beta=1/\alpha$  [lu2010zipfandheaprelation, baeza-yates_block_1997].
We focus our findings around Zipf’s law and Heaps’ law, as word counts are commonly used in natural language processing (NLP). We make the assumption that text will follow the form $N(t)\sim t^{\beta}$, evidence of which we describe next. While goodness of fit of this model and the underlying power-law distribution are often the subject of statistical scaling law research, our focus is on linguistic explanation; thus, we de-focus on goodness of fit.
Heaps’ law research has focused on written language heavily based in literary works or corpora [chacoma_tags_2020, bian2016scaling_like_lin, altmann2024statisticallaws, piantadosi2014zipf_mechs_review], such as Google’s N-gram corpus derived from books [gerlach_vocabgrowth2013].
Some of the literary work has gone further to examine scaling by parts of speech (nouns and verbs) [chacoma_tags_2020, piantadosi2014zipf_mechs_review] and by examining lemmas (root word forms) [corral2015lemmas_zipf], confirming good fit of both linguistic units as is often found with word unigrams.
The relationship between unique vocabulary and frequency is robust across many languages [piantadosi2014zipf_mechs_review, yu2018zipf_50langs_structural_pattern]; however, some work suggests exceptions to this robustness. Research using the Google N-gram corpus compared 7 languages’ more common lexicons and rarer lexicons, finding that in languages including Hebrew, Chinese, and Russian, rarer words are not Zipfian [petersen2012langs_cool_as_they_expand]. While finding good fit of lemmas in 4 languages, correlations between Zipf’s exponents for words and lemmas were not as robust for all languages, which could result from languages’ morphological complexity (English versus Finnish) [corral2015lemmas_zipf]. The type-token ratio divides vocabulary size (unique types) by total words to reflect text diversity; further work suggests that languages’ higher type-token ratio relates to higher morphological complexity and likely to produce many word forms [gutierrez_vasques2019morph_complexity]. Related, Zipfian and Heaps’ scaling do not hold for languages with limited dictionaries (e.g., Chinese, Japanese, and Korean) compared to English [lu2013zipf_lang_deviation]. These findings suggest that word segmentation and meaning encoding vary in languages and has scaling implications for languages producing many words. Thus, language and definition of a “word” (word-form) could drive steepness in scaling. [^1] Consider the morphological spectrum of languages from analytic to polysynthetic: polysynthetic languages generate novel word-forms continually (synthesizing concepts into one word), compared to analytic languages which tend to have fewer forms per word and relies on accompanying parts of speech for context [rijkhoff2002verbs].
Outside of classical literature, some textual analysis has looked at Wikipedia pages [altmann2024statisticallaws], internet forums [kubo_2011], news across languages [yu2018zipf_50langs_structural_pattern, rosillo_rodes2025entropy_ttr_gigaword], and Twitter [rosillo_rodes2025entropy_ttr_gigaword] to replicate Zipf’s or Heaps’ law. In online communities, while niche words marginally add to the lexicon, they can be core to a subset of the lexicon depending on the context [altmann2011niche_lexicons_online]—such as community-specific vocabulary in hobby communities—suggesting that context, if not medium, shows rich communication differences from a language’s broader, written lexicon. Yet little work seeks to study the mechanisms underlying scaling [piantadosi2014zipf_mechs_review] and language domains outside of written works.
Far less scaling research has focused on conversation corpora, and even less has offered social linguistic explanations of scaling mechanisms. The following small body relevant to conversation has been more varied than the research focusing on books or derivations. A notable work in this niche studied multilingual naturalistic corpora in mixed settings (face-to-face, telephone, and task-based), looking at multiple linguistic units of interest to include parts of speech and open and closed classes [linders2023zipf_spoken_dialog]. Ref. [linders2023zipf_spoken_dialog] went further to describe the scaling mechanism in spoken dialogue as stemming from the principle of least effort. A few other studies focus on spoken language transcripts. Research on elderly group conversation [abe_2021] studied whether cognition scores related to scaling exponents, finding cognition score groups showed some variation in Heaps’ law scaling exponents. Other work analyzed 10 transcribed naturalistic conversations from the British National Corpus and compared them to 10 books [bian2016scaling_like_lin]. [^2] Ref. [bian2016scaling_like_lin] seems to be the published version of Ref. [lin2015scalinglawshumanspeech]. Above are the most relevant studies to the current work, yet others have explored linguistic scaling laws in conversational data but with a different focus [baixeries2013children_adults_convos, torre2019buckeye_corpus_zipf, hernandez-fernandez_linguistic_2019, neophytou2017aphasia_zipf, ellis2009esl_convo_zipf]. [^3] The earliest relevant work here [ellis2009esl_convo_zipf] studied English language acquisition and assessed type-token scaling among learners. In a small sample of non-native and native speaker transcripts across very diverse settings, learners used highly frequent and prototypical verb constructions first—following Zipf’s law.
There is a gap in the comparison between naturalistic and monologue-styled text to explain underlying mechanisms for scaling. The monologue counterpart has usually been literary works focusing on unigrams of words as the informative unit of study [linders2023zipf_spoken_dialog], with little research beyond words [chacoma_tags_2020, linders2023zipf_spoken_dialog, yu2018zipf_50langs_structural_pattern]. It is worth exploring naturalistic conversation with respect to a non-naturalistic conversational counterpart because we would expect differences underlying these conversations beyond those of literary works (see again Table 1). These differences could help us further understand the uniqueness of naturalistic conversation—can we identify what makes spoken conversation naturalistic and other forms less so? It is necessary to expand the study beyond scaling regime of word frequency to explain these differences and what makes scaling meaningful in conversation. Thus, we see our work filling a gap in qualifying why there may be differences in conversational medium. Next we describe how linguistic theory can tie into scaling to understand these differences by medium.
1.3 Linguistic theory
Language explanations through information theory [shannon_mathematical_1948] and through content and frequency [estoup1916freqinshorthand] are some of the earliest examples of seeking to understand the mathematical relationships between the content of the language and the symbols used within it. Related work has led to studying efficiency in language such that word forms’ frequencies and later references balance competing pressures in language for clarity (descriptive enough) and brevity (without excess content). Numerous domains interrogate the concept of efficiency shaping language and how it appears in measurement. Zipf described the principle of least effort [zipf1949least_effort] to universally describe linguistic scaling. Work in information efficiency and psycholinguistics suggests it shapes the structure of language [futrell2024linguistic_info_efficiency, gibson2019efficiency_shapes_lang, zimmerman_tokens_2024, petersen2012langs_cool_as_they_expand] where human preference for communicative efficiency shapes language choices [fedzechkina2012restructure_lang_efficiency, jaeger2011efficiency_shape_lang_product]. Classical studies have found that text translations change and become simpler and worn down over time [west1973lectio_difficilior_potior]. And in pragmatic linguistics, Grice described this efficient tendency as a maxim of quantity in conversation in which one practices communicative brevity [grice1989wayofwords].
Grice’s Maxims prescribe rules governing conversation [grice1989wayofwords] that we can explore in conjunction (rather than in isolation) with scaling laws. The maxims state that interlocutors use quality, quantity, relevance, and manner of speaking to shape their utterances. These conversation maxims suggest possible distributions within the conversation flow that connect to the dynamic mix of language, cognition, and behavior producing these outputs. In particular, discussion of the maxims can be productive in understanding why universality exists across conversation modes. The maxim of quantity suggests efficient communication (supportive of linguistic scaling [xu2024wordreuse]) to achieve understanding with others. In different conversation modes, if transcribed naturalistic speech and writing for films both follow Heaps’ law, the cause could be underlying commonality.
From such works, we derive that (1) one significant purpose of a lexicon is to encode novel ideas—people do this through re-use and re-combination of words [xu2024wordreuse] and (2) efficiency, sometimes invoked as the principle of least effort [zipf1949least_effort], shapes the structure of communication [ferrer_i_cancho2003least_effort] and more broadly language across scales, e.g., [futrell2024linguistic_info_efficiency, gibson2019efficiency_shapes_lang, zimmerman_tokens_2024, fedzechkina2012restructure_lang_efficiency, jaeger2011efficiency_shape_lang_product]. The combination of these ideas suggests that using novel words for novel ideas also serves a valuable compressive purpose. Such compression helps speakers process structural constraints like information locality [futrell2024linguistic_info_efficiency], thereby directly tying vocabulary to efficiency. The maxims articulate behaviorally how speakers tie language to efficiency: they balance the pressures to both minimize cognitive load and maximize communicative diversity to achieve listener comprehension [grice1989wayofwords, futrell2024linguistic_info_efficiency]. Conversation is an interesting setting for achieving this balance in which we can apply NLP tools to measure potential mechanisms. While conversations tend to exhibit decreasing new types in language as described by Heaps’ law, conversation participants must be introducing or dynamically constructing meaning for the conversation to continue. This effect may be analogous to language novelty showing slow downs in production balanced by the need to construct meaning under new circumstances [petersen2012langs_cool_as_they_expand]. Next, we describe multiple linguistic units of meaning in conversation—tools we use—to explain scaling through the lens of linguistic theory.
1.4 Language tools for meaning
In addition to words (unigrams), we investigate the scaling of parts of speech (POS) to better understand their roles in meaning and scaling production in text. Parts of speech serve different communicative purposes in any text. Nouns and verbs primarily convey content whether new or referential to content already mentioned. Function words provide grammatical structure and connection in language, signaling meaning through grammatical and functional interactions with other words. Examples include the, and, that, many, not, and with. Interjections are exclamatory and not as syntactically related to surrounding parts of speech [universaldependencies2021]. Studying POS provides further context for Heaps’ law findings, as POS are used at different rates in communication. We explore how POS usage varies by medium and whether they show an impact on scaling.
Parts of speech belong to either an ‘open’ or a ‘closed’ word class. The open class is distinguishable for usually containing content words and having new words added from POS such as nouns and verbs. The closed class is more fixed and distinguishable for containing function words and numerals [universaldependencies2021]. We expect open class POS to show more novel types and therefore higher scaling exponents than closed class POS because open POS are typically used to add new content to a language. Closed class function words (common vocabulary) define a language, so close classes’ scaling should be relatively fixed at a large enough corpus scale.
Here, we examine interjections as well for their importance in conversations. Parallel speech is an important feature of naturalistic conversation in which listeners interject without interrupting while the speaker is talking, using linguistic entities like backchannels (e.g., mhm, yeah) and interjections (e.g., wow, really, oh my god[cooney2024naturalturn]. Interjections are a social linguistic gesture typically in the form of word, a non-word (e.g., pffft), or a phrase that can be indicative of a mental state or an emotive attitude or reaction [ameka1992interjections], reinforcing moments in a story or mirroring a speaker’s emotion [cooney2024naturalturn]. In this way, they are a significant source of interlocutor feedback [boudin2024multimodal_feedback_in_convo], even in text form which may have some gestural meaning encoded. That is, spoken interjections may pair with physical gestures that we have come to intuit as emotive feedback when encountering specific interjections. The meaning of this linguistic entity may rely especially heavily on embodied cognition and sensory or experiential grounding given interjections’ hypothesized co-evolution with communicating mental state or emotion [cartmill2025gestural]. Words often categorized as interjections can play numerous roles: phatic (social function), conative (directing attention), and expressive (where expression can be cognitive and/or emotive), making it difficult to rely on classifications based on surface word-form alone. We would not expect interjections to appear frequently outside of transcribed, spontaneous conversation (or its realistic fictional imitation) given interjections’ described roles and the scripted authorship constraints discussed in Section 1.1.
2 Current work
We explore the presence and steepness of Heaps’ law in naturalistic conversations between strangers as well as in fictional movie conversations in view of the previously described linguistic theory. Movie conversations convey story or plot to engage viewers and offer a different and understudied conversational mode. We build on prior work by analyzing conversation medium and comparing naturalistic to non-naturalistic explicitly. We further build on prior work by offering descriptive and exploratory insight through other NLP measures and breakdowns to understand POS differences. Some exploration was fruitful in qualifying results and laying the groundwork for future NLP metrics or methods to apply in conversation and statistical scaling domains. We utilize descriptive statistics such as POS proportions and type-token ratio, a metric in lexical diversity that has been used to compare languages [gutierrez_vasques2019morph_complexity, rosillo_rodes2025entropy_ttr_gigaword] and is directly related to Heaps’ law as diversity drives vocabulary size. The type-token ratio offers additional quantification of differences by medium and parts of speech. Last, we explore some temporal word distributions and variation therein [goh2008burstiness_memory, altmann2009beyond_word_freq] to understand how interarrival times of types and POS may show up in scaling and proportion results. As relates to parts of speech, some work suggests function words may occur more regularly than random while content words may occur less frequently but in more rapid intervals like clusters [altmann2009beyond_word_freq]. A study in naturalistic conversation also showed verbal communication tends toward irregular and clustered arrival times compared to non-verbal gestures [abney2018verbal_nonverbal_bursts]. Temporal and scaling analysis together may have implications for the two mediums in our investigation which we have not seen in our literature search.
Broadly, we expect differences between the naturalistic and non-naturalistic corpora in terms of Heaps’ law and descriptive statistics, offering a potential normative look into naturalistic conversation. Exploratory analyses contribute to our understanding of scaling differences between conversation mediums. We expect parts of speech scaling and usage to differ between open and closed word classes. We expect interjections to appear more frequently in naturalistic conversation.
3 Methodology
3.1 Data sets
We use the CANDOR dataset [reece2023candor] and the Movie-Dialogs Corpus [danescu-niculescu-mizil2011ChameleonsImaginedConversations] for the present project. We reference two formulations of data from the Movie-Dialogs Corpus: “Movies (individual)” as meaning individual conversations among speakers whereas “Movies (grouped)” means all conversations in an entire movie have been joined to represent a script that only contains the movie’s speaking parts; we refer to these corpora as Movies (I) and Movies (G) for brevity. Movies (grouped) reflects a movie’s plot or story as conveyed via dialogue only. We provide dataset description and specific cleaning and subsetting methods in each corpus subsection. We distinguish between ‘word’ and ‘type’ when counts refer specifically to words versus inclusive of all tokenized ‘types’ such as punctuation.
3.1.1 CANDOR corpus
CANDOR data was collected in 2020 [reece2023candor]. The corpus contains $1,456$ unique participants. All participants were adults of ages 18 to 66, and some participants completed more than one conversation (see Appendix A1 for more details). Demographic data by participant included age, sex, race, education, and employment status. We remove one conversation because the raw data showed it was shortened because of technical difficulties. The final count of unique participants is $1,455$ ($1,431$ with demographic data). The conversation data contains $1,655$ English language conversations, $315,938$ conversation turns (utterances), and 9.29.2M word tokens.
The conversation data exists pre-parsed into utterances using four different turn segmentation models: Audiophile, Cliffhanger, Backbiter [reece2023candor], and NaturalTurn [cooney2024naturalturn]. The choice of turn segmentation model will change the number of utterances, pauses, gaps, and counts of other between-turn variables. This project uses the NaturalTurn model which joins a speaker’s intended utterances while preserving a listener’s parallel speech as separate utterances, resulting in fewer utterances than prior segmentation models applied to CANDOR.
3.1.2 Movie-Dialogs Corpus: Movies (I) and Movies (G)
The Movie-Dialogs Corpus [danescu-niculescu-mizil2011ChameleonsImaginedConversations] contains fictional conversations diarized into ordered utterances from movie scripts. Release years range from 1927 to 2010 with movies heavily skewed to the 2000s. The dataset’s files contain separated metadata for movie titles, characters, conversations, and specific lines (utterances). Using these files, we reconstruct conversations (which are a list of roughly ordered utterance identifiers containing character identifiers). Data exploration shows inclusion of conversations with less than $10$ utterances, $2$ non-fiction genres, and $274$ null utterances fields spanning $166$ conversations. The current research questions pertain to content conveyed through conversation, so we remove conversations containing null fields as they could introduce erroneous signal. We subset the data to represent longer length fictional conversations and to enable more robust statistical analysis. Thus, we remove conversations with these criteria: having less than 10 utterances in the conversation, belonging to non-fiction genres, and containing null utterances.
The final corpus contains $2,397$ unique characters from $472$ unique movies throughout the release years 1927 to 2010. The conversation data includes $3,104$ conversations, $43,762$ utterances, and $485,389$ words. This corpus forms our Movies (individual) comparison group.
Movie conversations are very short compared to CANDOR’s naturalistic ones (see Table 2), so we formulate another exploration of Movie scaling to look at all conversations within one movie as a ‘conversation.’ Movie-Dialogs data does not ensure chronological ordering of all conversations within a movie, but prior work shows scaling replicates when compared to shuffled bodies of its own respective text [altmann2024statisticallaws]. The conjoined speaking parts formulation as another ‘conversation’ is conveying plot solely through dialogue from numerous participants. Still excluding conversations with null utterances or movies in non-fiction genres, this formulation allows us to include $8,481$ unique characters, $589$ movies (as $589$ conversations), totaling $288,125$ utterances and 33M words, and to have another comparison group called Movies (grouped).
3.2 Tools
All coding occurs in a Python 3.12 environment in Visual Studio Code, with libraries including NumPy, Pandas, and Matplotlib. See the project repository for dependencies and versions. Compute resources include an Apple M1 Macbook Pro with 32 GB of memory. Statistical analyses rely on SciPy for linear regression. Natural Language Processing (NLP) relies on regular expressions (re library) and NTLK for tokenization while Stanza’s pretrained English model [qi_stanza_2020] provides part-of-speech (POS) classification. Stanza’s model was trained on treebanks (containing linguistic dependency information), and Stanz’s POS processor utilizes sentence context in tag classification. Token labels come from universal part-of-speech tags [universaldependencies2021]. Stanza’s model shows high accuracy averages. [^4] Summarized from Stanza’s published model performance by language and treebank where the utilized treebanks are highly relevant to the current discourse types.
3.3 Procedure
3.3.1 Data cleaning
We requested and received CANDOR data from its respective authors [cooney2024naturalturn] and downloaded the Movie-Dialogs dataset in its legacy format [movie_dialogs_data_legacy_url]. [^5] The Movie-Dialogs dataset is available from multiple sources, including the ConvoKit Python package.
Data cleaning includes various NLP steps depending on the conversation mode to bring the written artifacts in movie scripts more in line with spoken conversation in CANDOR. For CANDOR, we remove a transcription artifact in NaturalTurn ($\ll$’ and $\gg$’). For movies, we remove some grammatical artifacts such as standardizing quotes and hyphens, and we fix some encoding issues and potential whitespace issues. For all data, we spell out shortened word-forms: sorta, dunno, gonna, wanna, gotta. The last NLP step is POS tagging using Stanza’s processor per corpus after which we recode POS tags to fewer categories (see Appendix A3). For one of our exploratory variables, we code conversations as having low or high interjection usage using a median cut within corpus.
After some comparison and careful conceptual consideration of punctuation’s place in the text form of conversation (it may convey prosodic information), we tokenize the cleaned text with punctuation kept in. Punctuation can convey meaning and has statistical importance in showing Zipf’s law [stanisz2024nlp_complexity]. Removing punctuation mildly influences POS tagging and subtly changes scaling coefficients while increasing standard error of the slopes. For instance, interjection (INTJ) occurrences are 1.5M without punctuation and shrink to 598k with punctuation because punctuation acts as context to bias INTJs downward; it is possible that punctuation removes INTJs that are spoken as utterances and could serve a continuer purpose (such as “ok.” spoken while listening).
3.3.2 Statistical procedures
To explore the data, we use Pearson correlations, observe distributions, and analyze descriptive statistics. We also explore temporal measures of words and POS tags such as token interarrival times, burstiness, and memory [altmann2009beyond_word_freq, goh2008burstiness_memory]. Interarrival times are gaps between successive occurrences of types. Burstiness codifies the variation of interarrival times in a metric ($1<\text{B}<1$), with $\text{B}>0$ suggesting ‘bursty’ behavior (less regular) behavior while $\text{B}<0$ indicates more regular behavior. Memory is a correlation measure of interarrival times that codifies if these gaps are independent or show memory ($1<\text{M}<1$). $\text{M}>0$ indicates that similar length gaps follow each other (short and short or long and long). $\text{M}<0$ suggest more alternation between short and long gaps. Last, we calculate the type-token ratio (TTR) as a measure of lexical diversity or richness [gutierrez_vasques2019morph_complexity, rosillo_rodes2025entropy_ttr_gigaword] as vocabulary size (unique types) divided by total words where a value closer to 1 reflects high diversity.
We assess Heaps’ law and Zipf’s law with least-squares linear regressions, visually identifying the upper scaling regime for Heaps’ scaling behavior. We collect total and unique types over the length of the document. We visually identify a restricted range of the data to isolate the sublinear scaling regime in log-log space. The heavy tail of frequency distributions presents some measurement difficulty, one of which is violating the normality assumption underlying least squares regression. Restricting the regime allows us to still use regression to study the steepness of slopes [clauset2009powerlaw_regimes].
We apply a horizontal axis restriction in log-space for the Zipf regressions (ranking of types) and correspondingly apply those same restrictions to Heaps’ regressions’ vertical axis (unique types). These restrictions bring the expected correspondence between Heaps’ and Zipfian findings better in line. Given the shortness of data in Movies (I) and some POS, we could not apply the same end restriction to all analyses (they would be null if there were not sufficient data). We design start and end-ranges matrices to specify values per analysis (in Appendix A3). Further, given that the plot lines for regressions reflect unique types over the totality of a corpus or category, length of documents on average makes some plots’ lines appear longer or shorter; in particular, Movies (I) has much shorter documents, leading to their total ‘types’ (plot line) falling short of the other corpora.
Next, we run a least-squares linear regression per corpus, POS by corpus, and low versus high interjection usage by corpus.
4 Results
4.1 Descriptive analyses
CANDOR Movies (I) Movies (G)
Utterances $191$ $(66.42)$ $14$ $(5.56)$ $489$ $(221.55)$
Words $5563$ $(1556.26)$ $156$ $(92.49)$ $5122$ $(2463.41)$
Speaker words $6327$ $(5459.16)$ $202$ $(268.94)$ $356$ $(560.50)$
Table 2: Raw descriptive statistics of utterances and words by corpus and words by speaker. Values are reported as: $\mu$ ($\sigma$). See Appendix A1 for word by utterance relationships in conversations and more spread statistics.
[…]
Figure 2: Words by utterances heatmap per corpus with Pearson correlation coefficients. The Movies (individual) corpus is subsetted to include conversations having a minimum of 10 utterances.
Corpora descriptive statistics ($\mu,~\sigma$) for number of words and utterances at the conversation level can be found in Table 2. CANDOR and Movies (G) result in the most similar structure for number of words while Movies (I) is distinctly much lower for both categories. Given the wide-ranging variance, we look at coefficients of variation ($\sigma/\mu$) to quantify the spread further. Of the six categories (words and utterances per corpus), number of words for Movies (I) ($0.59$) and Movies (G) ($0.48$) show the highest coefficients of variation (compared to $\text{CANDOR}=0.28$). Related, we look at the type-token ratio (TTR) and average across conversations’ TTRs at the corpus level. The corpus TTR distributions are normally distributed, and we report TTR as $\mu~(\sigma)$. Each corpus’s ratio appears distinct, with CANDOR (0.17(0.02)0.17~(0.02)) having the lowest diversity, followed by Movies (G) (0.24(0.07)0.24~(0.07)) and Movies (I) (0.64(0.09)0.64~(0.09)) with the highest diversity.
Further, the relationship between number of words and number of utterances is much stronger in Movies (grouped: $r=0.89$, individual: $r=0.67$) than in CANDOR ($r=0.32$). Movie text appears less consistent and more subjective to outliers than CANDOR. Overall, Movies (G) is distinctly linear (as seen in Figure 2).
[…]
Figure 3: Normalized part of speech corpus proportions as total words (A) and unique words (B). A corpus’ total and unique proportions each sum to $1.0$ within corpus. Any likeness to country flag is coincidental in our endeavor to use accessible color scheming.
Next, we analyze conversations’ unique word stagnation to detect outliers. A ‘run’ of unique words indicates stagnation in the arrival of new words. We first identify that the distributions of conversations’ maximum runs are right (positively) skewed toward shorter runs. CANDOR ($55$) and Movies (G) ($48$) show very similar run medians with more variance for movie scripts. Movies (I) ($6$) skews the shortest. We identify the run value at which $99.95\%$ of data is included in the distribution and apply this run value as a cut-off to detect outliers. Figure A7 highlights these outliers. CANDOR’s outlier results in one of the corpus’s longest conversations. Movies (I)’s outliers include a singing exchange from The Rocky Horror Picture Show and a scene from The Leopard Man; we note that The Leopard Man’s result identifies a duplicate scene included in this conversation due to it being an alternate scene that may have been accidentally scraped. The Rocky Horror Picture Show exchange shows the same lines being exchanged between the character Rocky and guests, resulting in low novelty. Movies (G)’s outlier is the dramatic film, Magnolia.
Parts of speech ($\mu,~\sigma$) are in Table A2 where averages are reported as total words and as unique words in each corpus. POS proportions per corpus are in Figure 3. Despite Movies (G) often having the highest variance across POS, CANDOR and Movies (G) appear more similar in some categories than does Movies (G) to Movies (I).
[…]
Figure 4: Zipf’s law (A) and Heaps’ law (B) for the corpora. The grey dashed line indicates (coefficient =1=1), and the solid black line indicates the scaling regime restriction on the (A) horizontal axis or (B) vertical axis. Standard error of the slope: no notation for SES <.01<.01, ˆ for $05>\text{SES}\geq.01$, and $\ast$ for SES >.05>.05.
Looking at proportion of POS usage by corpus, ‘other’, noun, and verb in order are the largest categories for total usage while noun and verb are the largest categories for unique usage. For unique noun proportions, Movies (G) ($0.39$) is higher than CANDOR ($0.32$) and Movies (I) ($0.29$) (see Table A3 for top occurring interjections). Further differences are more apparent in function words and interjections: Movies (I) has the most unique function word usage ($0.09$ versus CANDOR =0.06=0.06 and Movies (G) =0.04=0.04). CANDOR shows the most total interjection usage, but this proportion shrinks in unique usage ($0.11\rightarrow 0.05$); Movies (I) and Movies (G) interjection usage remains at .03.03 for total and unique proportions. When adding together our categories of ‘other’, function, and interjection POS, this aggregated POS proportion in CANDOR accounts for $0.58$ of total words (compared to $0.48$ in both Movies (I) and Movies (G)); CANDOR differs from Ref. [chacoma_tags_2020]’s ‘other’ proportion ($0.50$) while movies’ usage almost agrees. This difference results from CANDOR’s noun proportion ($0.22$) being lower than Ref. [chacoma_tags_2020] ($0.31$). Verbs are comparable across our sources to Ref. [chacoma_tags_2020].
Exploratory analyses inform our understanding of features and help to qualify the study of scaling. We reference some of these findings, with additional notes and results in the Appendices.
4.2 Temporal analyses
We analyze conversations’ narrative thirds (tertiles). For tertiles’ interarrival distributions, all corpora’s gaps are highly skewed short but with long tails (Figure A4). CANDOR and Movies (G) resemble each other in this regard, though CANDOR’s tail spans slightly more magnitudes in the beginning and middle of conversations. We summarize the extremely high interval tails by summing the tails’ word appearances as POS proportions by corpus (see Table A4). CANDOR’s and Movies (G)’s highest tail gaps are nouns and Movies (I)’s are ‘other’. Qualitatively, we see examples of punctuation and pronouns as some of the shortest intervals.
We split interarrival distributions by POS (Figure A5). POS interval differences become more apparent when comparing total types to unique types. Noun, verb, and ‘other’ POS total gaps are much shorter than function word and interjection usage (but with longer tails).
CANDOR has the most regular interjection usage, demonstrated by shorter gaps. Compared to other POS, interjections and function words show much longer wait times for new unique types, suggestive of the closed class definition of these POS. Unique types’ interquartile range reflects extremely long intervals before new types appearing, capturing Heaps’ law behavior.
We calculate burstiness (B) and memory (M) along with multiple other comparison groups as we have not seen many textual results in this dual-metric format, excepting Ref. [goh2008burstiness_memory]. In Figure A6, we include B and M for: corpora, POS, literary works as a comparison (Frankenstein, Pride & Prejudice, and Moby Dick), and shuffled versions of each body denoted with $\ast$. We find differences for these literary texts compared to the 2 books reported in Ref. [goh2008burstiness_memory] which has burstiness and memory of effectively 0. Our literary comparisons cluster around $\text{B}=-0.09$, $\text{M}=0.35$. CANDOR and Movies (G) are close in burstiness to this literary cluster but lower in memory. Movies (I) (and its POS) shows the least memory (close to 0) but the most regularity (negative burstiness). Parts of speech differ more starkly from CANDOR and Movies (G) averages. Their interjections and function words are close to $\text{B}=0$ and $\text{M}=0$. The rest of the POS cluster around $\text{M}=0$ but trail into the lower left quadrant of less memory and less burstiness, which suggests more alternation between interarrival times and more regular behavior than random. Notably, all the Movies (I) POS are in the lower left quadrant. See Appendix A2 for temporal figures.
4.3 Regression analyses
We observe sublinear scaling and Heaps’ law similarly to prior works, e.g., [chacoma_tags_2020, linders2023zipf_spoken_dialog]. CANDOR reliably has longer conversations, and scaling exponents for individual conversations are normally distributed (Figure A1). For Movies (individual), conversations are less consistent and short. Each of the three corpora show CANDOR: $\beta=0.65\pm 0.0001$, Movies (I): $\beta=0.64\pm 0.003$, and Movies (G): $\beta=0.63\pm 0.0003$ (see Figure 4 for Heaps’ and corresponding rank regressions).
Following the same process, we examine POS by category by calculating linear regressions per POS per corpus on restricted scaling regimes (see Figure 6). Movies (individual) shows the highest standard errors of the slopes given their shorter and more variable length conversations. We note much higher standard error of the slope for Movies (I) when we keep punctuation in the analysis, which may be due to scaling regime changes from pre-punctuation (i.e., sensitivity to regime definition). Nouns’ coefficients are high across corpora (CANDOR being the highest) and higher than the corpora’s overall $\beta$ values. Every other POS shows stark differences by corpora. Verb and function word scaling are highest for Movies (I). CANDOR shows the highest coefficient for ‘other’ POS. Interjections are highest in Movies (G). This result is a post-punctuation change from interjections being the lowest in CANDOR (now second) compared to Movies (I) (now last). We note the lower number of total interjections and some difficulty in fitting.
Last, we regress on low and high interjection categories. The group proportions for low and high interjection variable usage are: CANDOR =[0.51,0.49]=[0.51,0.49], Movies (I) =[0.60,0.40]=[0.60,0.40], and Movies (G) =[0.50,0.50]=[0.50,0.50]. Split interjection analyses show only slight variation from their average corpora coefficients and by group. With [low $\beta$ and high $\beta$]: CANDOR =[0.68,0.67]=[0.68,0.67], Movies (I) =[0.75,0.72]=[0.75,0.72], and Movies (G) =[0.68,0.65]=[0.68,0.65].
[…]
Figure 5: Part of speech linear regressions for the corpora. Below the grey dashed line ($\beta=1$) is sublinear scaling, and the solid black line indicates the scaling regime restriction on the vertical axis. Standard error of the slope: no notation for SES <.01<.01, ˆ for $05>\text{SES}\geq.01$, and $\ast$ for SES >.05>.05.
[…]
Figure 6: Part of speech rank regressions for the corpora. The grey dashed line indicates $\alpha=-1.0$. The solid black line indicates the scaling regime restriction on the horizontal axis. Standard error of the slope: no notation for SES <.01<.01, ˆ for $05>\text{SES}\geq.01$, and $\ast$ for SES >.05>.05.
5 Discussion
5.1 Presence of Heaps’ law in conversation
Our work identifies Heaps’ law scaling in naturalistic and fictional conversations. We find that scaling exponents are normally distributed for naturalistic conversations, consistent with multi-participant internet forums [kubo_2011]) and with past scaling behavior for spoken settings [abe_2021, bian2016scaling_like_lin]. The normal distribution of scaling exponents intriguingly offers a statistical artifact for future conversation research to replicate or to contrast to other forms of conversation or text. When comparing the three corpora, naturalistic and conjoined movie script dialogue are the most statistically similar across total words and type-token ratio, with both showing much lower variance around Heaps’ law exponents.
That the trajectory of a movie plot (as communicated through conversations) has diminishing vocabulary additions in common with naturalistic conversations suggests that the same information efficiency mechanisms underlie both. Of course, many of the same rules shape the language used in both contexts; for example, common words will be introduced early and repeated, inevitably leading to a diminishing innovation rate. The scaling commonality among both contexts is surprising more so because of the contexts’ differences in authorship and purpose. Movie scripts are often collaboratively written by fairly large teams with disparate artistic visions and incentives over extended periods of time. Other research on multi-participant discussions also follows Heaps’ scaling behavior, both in internet forums [kubo_2011] and group discussion [abe_2021] but suggests that having more participants could contribute to novel word production. However, we do not see higher novelty production in movies than in naturalistic conversation. The adherence of movies to scaling similar in CANDOR is interesting because of Movies (I)’s distinctively low number of words and utterances but very high lexical diversity (a much higher type-token ratio than the other two corpora). The lexical diversity makes sense for individual movie conversations, as: 1) they are designed to communicate plot points and need to be differentiating per plot point to do so, and 2) plot points are likely variable per movie because of different artistic involvement. Movies (I)’s necessary lexical diversity, while following scaling behavior at the corpus level, could be why we see such high variance in their scaling exponents as well as in number of words and utterances.
We further delve into this lexical diversity. Despite their shortness and sparsity, single movie conversations are imbued with layers of multimodal content (e.g., cinematography, music) to help the audience make sense of the plot as they experience it over the course of the movie. These short conversations can draw on the broader scaffolding intent of the entire movie, further pointing to a role that individual conversations serve within their larger structure. Meanwhile, naturalistic speech has different layers of content—gestural and prosodic cues—with spontaneous sense-making. Movies imbue the intended perception at the planned times and build on themselves to convey holistically bigger themes in the movie (e.g., coming of age, justice) that the short dialogues alone may not convey because of content sparsity. [^6] Reminiscent of the Aristotelian concept ‘the whole is more than the sum of its parts’ and Gestalt theory, the effects achieved at the level of single conversations compared to entire scripts speak to possible emergence. Future research could ask: Are there computationally identifiable, emergent differences in how an entire movie achieves holistic themes not seen among the movie’s cumulative parts?
Overall, both fictional and naturalistic modalities rely on their respective layers for storytelling and conveying meaning. Presuming a conversation is conveying themes, even spontaneously, the design of movies to convey themes holistically (and its apparent data make-up) makes the entire conjoined script more analogous to a natural conversation than to a sequence of sparse, short conversations.
5.2 Part of speech differences by medium
Having identified global scaling behavior at the corpus level, we next focus on scaling behavior among parts of speech. We find open parts of speech to have higher scaling than closed parts of speech; the scaling exponents for nouns and verbs (open classes) are higher than other categories within the corpora—similar to behavior of nouns and verbs in literary works [chacoma_tags_2020], with our ‘other’ category having the next highest scaling. Nouns exhibit relatively higher scaling exponents across our datasets than do other POS, with nouns being higher than the average exponents for each corpus. Only verbs notably follow this higher-than-average pattern in Movies (I). This commonality to scaling in literary works [chacoma_tags_2020] suggests some universality in the way the English language informs listeners or audiences, with POS suggesting differences by medium. Higher production of nouns (CANDOR, Movies (G)) and verbs (Movies (I)) relates to structure and meaning-making. The former two texts have longer and more informative messages to convey. Individual movie conversations are very short and may rely on verbs, references to who is doing what or events taking place, in dialogue instances to be as informative as possible during their short-length occurrences [martin2010pos_bnc_movies]. Consider the proportions of unique nouns and unique verbs that appear: they appear the most across all corpora out of all the POS, with Movies (G) showing notably higher unique nouns. Conjoined movie conversations likely rely on nouns to encode novel plot content over the course of films beyond single movie conversations and naturalistic conversations. A conjoined movie conversation represents an entire plot through speaking parts. These speaking parts represent a complex causality, a phenomenon itself that is very different to spontaneous speech, in which the speech and world are co-created and intended to co-occur to progress the plot. An important factor in the context of conversation involves relation-management: interlocuters’ relation and how speakers relate to their stances [ruhlemann2007ConversationContext]. Nouns encode the arrival of new plot points, suggesting multiple pragmatic functions. In particular, proper nouns aid in relation management, offering a way to linguistically connect people through social and empathetic expressions [martin2009teaching_convo_film, ruhlemann2007ConversationContext]. Further, pronouns often co-occur in speech acts, which abound in film dialogue. Relevant speech acts could include expressing politeness, thanks, or apology, where the use of “I” is abundant [martin2010pos_bnc_movies]. Overall, these examples illustrate how the higher scaling of nouns suggests more active plot encoding or management of how character for the purpose of relating well to the audience.
In Movies (I), function word scaling is much higher than in the other two corpora. Higher function word scaling could reflect a preference for greater clarity or precision within shorter bouts of dialogue. This precision could appear as more complete context rather than relying on ellipsis, the omission of words that one can infer through context (finding “meaning without form”[merchant2018ellipsis_chapter]). [^7] The linguistic study of ellipsis is complex and contains numerous types. See the Oxford Handbook of Ellipsis for further study. Some examples of providing greater context in movie conversations could include 1) omitting actions but having characters talk about those actions or 2) character dialogue about events without dropping redundant words or phrases. One way this inclusion helps the audience could be to work around an information locality constraint such that characters reference much earlier events in sufficient detail to further the more current plot point and achieve audience comprehension—minimizing cognitive load [grice1989wayofwords, futrell2024linguistic_info_efficiency].
Notably, our investigation of burstiness and memory yields a similar result. Movies (I) consistently show the most regularity as to when all of this dataset’s parts of speech occur, a regularity associated with function words [altmann2009beyond_word_freq]. This finding bolsters the thorough construction explanation of movie conversations. The idea of greater content clarity through function words also follows from movie dialogue possessing a stable and repetitive lexico-grammatical profile [zago2020film_discourse]. The short nature of Movies (I) conversations and high function word usage suggests that function words help structure meaning for viewers.
We find that interjections’ scaling exponents are intermediate to those of open (noun, verb) and closed class (function words) scaling for the larger corpora (CANDOR and Movies (G)), which is consistent with interjections being a semi-open class [ameka1992interjections]. Movies (G) shows the highest scaling in interjections, but CANDOR shows the highest total and unique proportions of interjection use. This discrepancy between high scaling (type novelty) and usage (both new and repeating types recurring) could be a factor of document length, dual participants, and the nature of spontaneous conversation. Interjections serve an important feedback and reinforcement role in conversation between interlocutors [cooney2024naturalturn, boudin2024multimodal_feedback_in_convo] which is not functionally needed in prepared dialogue because it is more cohesively organized [martin2009teaching_convo_film, zago2020film_discourse], lending to lower use in movies [martin2010pos_bnc_movies]. It is apparent, at least, that the corpora rely on different interjections (Table A3), suggesting that different types of interjections are better suited to the purposes of conversation in each corpus; for instance, naturalistic speech shows more space-filling words (‘like’, ‘uh’, ‘um’) while we see more tags in fictional speech suggesting the introduction of concepts or contradiction (‘well’) and general use of ‘yes’ and ‘no’. Among the corpora, interjection usage is uniquely high in CANDOR, suggesting naturalistic conversation may rely on interjections as a connective supplement between participants and the meaning they attempt to convey (more-so than in movies [martin2010pos_bnc_movies]). In movies, actors are not trying to relate to each other in the same way as in spontaneous conversation (where interjections would be relational). Writing and physical movement are pre-scripted, so actors may not intuit or think to add vocal gestures despite attempting to act how humans act and sound. For the low and high interjection usage groups, we find no difference in group scaling within corpora. Interjections are an important part of building rapport, but they do not seem to have an additive effect on content in conversations. This null result suggests it may be more pro-social than informational, consistent with the conception of interjections as vocal gestures rather than speech [ameka1992interjections]. Another explanation is that linear regression is a limited method for delving into interjections. First, interjections are a difficult part of speech to parse automatically as its word-forms and recognizability highly varies. A POS tagger could miss instances, contributing to low counts and high uncertainty in a regression estimate. Second, interjections’ pro-social purpose provides more feedback to the listener than is contained in a single word expression, and analyzing unique and total words in a regression simply misses this context in many NLP analyses.
5.3 Conveying meaning
Last, we end on some thoughts and theory around how the affordances of each corpus help to convey meaning. Each corpus has a part of speech that sees notably high use; this usage suggests that the conversational medium contains more novel forms to encode new information rather than relying on old forms already seen. As we discuss prior (Section 1.3), lexical efficiency refers to the principle of least effort [zipf1949least_effort] where the preference to communicate efficiency shapes language and structure choices [fedzechkina2012restructure_lang_efficiency, jaeger2011efficiency_shape_lang_product, ferrer_i_cancho2003least_effort], such as re-using or combining words in different ways [xu2024wordreuse]. We suggest these higher-use categories are inefficient, offering high utility in communicating precisely in the modality in which they are more highly used. The lexically inefficient POS we refer to are nouns and verbs in Movies (G), function words in Movies (I), and interjections in CANDOR. These POS may be more central to interlocutors understanding what is being communicated in that medium.
One reason for this language inefficiency could be the tendency to provide the most relevant information according to the delivery method—whether in-person or a writer imagining an audience’s reaction; under the maxim of quantity [grice1989wayofwords], the speaker has to judge how to provide maximally useful content so the listener comprehends the intent. For speakers to deem content maximally useful, speakers must adapt the dialogue to the circumstances in which they deliver it, considering the person receiving it, what layers of activity they can control (for example, speech, gestures, or writing direct messages), time of day, and other events taking place—circumstances broadly covering the dimensions of time and space [rijkhoff2002verbs]. Adapting to these circumstances also fits with Marconi’s two-strategy lexical competence explanation wherein inferential competence relates to structure and referential competence relates words to real entities [marconi1997LexicalCompetence] or representations thereof [varnier2000DiegoMarconiLexical]. The reliance of each corpus on particular strategies lets us consider how each corpus’s mean-making mechanisms may differ compared to standard speech. The different modalities’ inefficient parts of speech present a sort of fingerprint to make the most understanding in what is being conveyed, leaning more heavily on either lexical or referential competence; naturalistic and entire movie scripts’ reliance on content words (nouns and verbs) provides more pointers to the world and concepts. Single movie conversations’ reliance on verbs also makes use of this referential strategy but leans more heavily on lexical competence than the other corpora through function words. Speakers deliver naturalistic conversation using spoken utterances without a pre-determined structure to turns, allowing them the benefit of listeners’ interjections and other backchannels that helps speakers determine whether the listener understands the message. Movies in their entirety have the benefit of time and multimodality to convey their message. Instances of movie conversations have the challenge of ferrying audience from diegetic point to point, requiring that they maximize the space (perhaps the conversational floor) via more structure and specific verbiage. Actors do not have the benefit of interjection feedback from listeners to know the message is being perceived well.
Overall, speakers or writers try to infer others’ mental states to communicate with them effectively [astington2005theory_of_mind_language]. With different affordances for communication, mediums shape communicative behavior via how speakers connect and know they are connecting.
Speakers can leverage Theory of Mind and can maximize the space the conversational medium offers. Conversation itself is a communication tool in which we semantically encode and present our mental states, sometimes to convey complex things to other people [astington2005theory_of_mind_language]. Although much signal from prosody flattens in transcribed text, our work’s interjection results in transcribed spoken conversation suggests some mutual information with what is happening during spontaneous conversation [wolf2023prosody_text_redundancy]. In spontaneous conversation, speakers control the layers of activity of gestures and expression to convey their intended meaning. Interjections are both vocal gestures and, other times, expressions that enable listeners to show they have both perceived and understood the message—listeners are using theory of mind to demonstrate their understanding of the speaker. Authors of movie scripts access these layers but can also control numerous other elements to support communicative intent to help ‘listeners‘ understand the mind of the other. These elements could be lighting, sound, or costume to show emotions or state of mind. Writers in particular use mental models of people, situations, concepts, and etc. to represent concepts within speech. However, these mental models are communicated one way (without feedback and amendment to the message).
5.4 Limitations
The parts of speech classification method we apply relies on linguistic dependency trees (treebanks). The stanza model for POS tagging will err in the direction of context from trees rather than classification based on surface-form (e.g., lexicon). Researchers could compare stanza output to other POS processors such as NLTK, spaCy, or large language model tagging. Differences in tagging models would affect conclusions drawn about forms of text conversation (e.g., spoken and discussion platform).
Alternate natural language processing (NLP) decisions would change our results. In all data cleaning procedures, we maintain as much of the original data as feasible, and we run parallel analyses on more- or less-cleaned data to inform our choices, which impact the distributions of token, part of speech, and interjection frequencies. We run our analyses with and without punctuation, ultimately relying of analyses with punctuation; punctuation can convey content from prosody [wolf2023prosody_text_redundancy], and including it with the rest of text can more completely express Zipf’s law [stanisz2024nlp_complexity]. When important, we note throughout our work the substantive differences we find between these two instances of the data (without versus with punctuation). Another possible limitation arises from cleaning the mediums to be more comparable, such as the removal of grammatical artifacts from fictional dialogues under the assumption that these would not be acted or spoken. These NLP choices could exclude the unique signal found in movie conversations. A future project could investigate whether there are meaningful effects of NLP and medium on the fit of Heaps’ law by comparing additional data cleaning pipelines applied to the same corpora. Likewise, our choice of conversation segmentation algorithm shapes the data. Research could compare conversation segmentation algorithms to see how they might change the shape of conversations.
From the perspective of information theory, some may interpret the purpose of Heaps’ law as an information metric. However, the technical formalization of ‘information’ as Shannon entropy does not typically equate to meaning [shannon_mathematical_1948]. We interpret differences among scaling law coefficients as a tentative window and way to narrow our focus into explanatory mechanisms; a higher coefficient for one corpus does not mean that corpus conveys ‘more’ meaning, for instance. Thus, one limitation of the current work is the degree to which we can assign meaning to changes in scaling regime. We suggest future work to bolster explanation by continuing to focus on linguistic mechanisms of scaling [piantadosi2014zipf_mechs_review].
Another limitation comes from restricting a regression’s scaling regime to a subset of data using a visual method. This method is sensitive to the selected range (choice of $x_{\text{min}})$. If researchers’ goals involve fitting data to the power law distribution explicitly and evaluating goodness of fit, a more expansive method may be desired [clauset2009powerlaw_regimes] that does not require committing to an a priori data range.
5.5 Future work
Building on our current work, we discuss avenues for future applications in film analysis, technology medium, and conversation. First in film analysis, research could adapt the Movie-Dialogs corpus’ film selection methodology and the corpus’s skew toward short conversations. Over three decades, social media text length consistently decreased and communication norms changed [di_marco2024linguistic_simplification], suggesting the same could be true in scriptwriting. Research into the evolution of fictional storytelling could: Are movie conversations globally short? How did storytelling evolve to become closer to the kind of storytelling we experience circa 2025? How has script construction changed? Have themes in fictional storytelling substantially shifted over decades? Future work could also bolster movie title sampling by: adding movies with longer conversations; sampling titles according to selection criteria (e.g., ‘top of’ list per year such as grossing, rated, or liked); or identifying a release year after which the evolution in movie storytelling ‘stabilizes’ such that there is some computational cohesiveness to the subset of movies—from which we can say, for example, that these movies are of a type or that they contain the storytelling or script construction more common to the last three decades of movies. Second, film analysis would benefit from transcribing dialogue as performed to capture deviations from official scripts by recording performed pauses and any naturalistic features injected by actors (backchannels and interjections). Another avenue is to code these linguistic features as a ‘type’ in the text and analyze scaling behavior with them included; this approach begins to answer the role of pauses and interjections in cognitive processing.
One could consider the current work’s mediums alongside literary works, texting, and online communication to better situate conversation data from naturalistic to non-naturalistic. Naturalistic speech has a more instinctive construction than technological communication by virtue of the ability to prepare and edit discourse via the latter. We should ask what it means for communication and cognition when dialogue is spontaneous versus pre-planned; how do mediums’ affordances cognitively and behaviorally drive conversation? We compare scripted to naturalistic conversation, yet much communication occurs technologically. Though non-naturalistic, online discussion forums share more qualities with naturalistic conversations than do scripted conversations. For example, Reddit users develop a causal, conversational posting style. [^8] On Reddit, posters use orthography to convey prosody and incorporate internet- or community-specific speech to sound more like speech. A post could start with demographic information, e.g.: “I (24f) and my boyfriend (23m)”. Unlike spoken conversation, Reddit’s structure has an enduring social feedback mechanism where other users can encounter and judge posts. Discussion forum posts may receive feedback through the medium and see author edits to the original language, making the language susceptible to co-construction at the platform level. Qualities of the discussion forum include turn delays and cognitive preparation. This medium features symbolic supplemental communication including emojis and gifs. Data wrangling considerations include possible post deletion by users. Future work could quantify how mediums supplement flattened information to imbue meaning in a toneless medium.
Conversation research can gain further meaning from text. First, conversation topics would be fruitful in identifying meaning. Regression coefficients reflecting Heaps’ law scaling and cosine similarity are simple ways to start exploring how to identify abrupt and meaningful topic shifts and callbacks across conversation windows. These minimalist topic markers could provide context and be compared or used in conjunction with models such as BERTopic. Second, interjections offer rich signal that could be furhter disambiguated by type. A common type distinction is interjection words versus interjectional phrases, often thought of as parts of speech versus syntax. Interjections whose primary role is feedback can exhibit uncommon letter sequences and sounds (e.g., pfft, tut-tut[ameka1992interjections]; paired with the syntactic construction of interjectional phrases, interjection forms have tagging and token implications and may show different scaling between types when correctly recognized. Interlocutor feedback takes various forms—including the lexical form on which we focus—and fulfills many roles, such as that of continuers, expressing a reaction, and acknowledgment. Consistent with this purpose, we find interjections’ appearance cadence in naturalistic conversation to be more regular than in scripted conversation, and we find more use of typical filler words in naturalistic speech. Generic feedback, such as interjection words, likely occur during conversational set-up, whereas specific feedback (such as interjectional phrases) is more closely connected to the local context [boudin2024multimodal_feedback_in_convo]. Future research could pair the timing of interjections with their word-forms to qualify when the purpose of their usage changes (such as simply showing interest, good manner or showing common ground).
Third, codifying the motivation for conversation may reveal what we can expect from different conversation avenues. Each avenue may convey a different conversation purpose, from Movies and CANDOR to other corpora like NPR’s empathy-laden StoryCorps,999storycorps.org/discover/archive the Structured Podcast Research Corpus (SPoRC),101010github.com/blitt2018/SPoRC_data or the BabyCenter forum. Podcasts alone have relatively recently risen in popularity and host many genres or purpose-driven series (such as entertainment, education, news, politics, and self-improvement).
Lastly, the nature of conversation provides some broad avenues for Heaps’ law applications. Questions include: how quickly do distributed versus in-person conversations stabilize and do controversial conversations have different markers? These research questions have implications for interlocutors reaching shared understanding, seeing fruitful discussion on platforms, or distinguishing people from bots.
6 Concluding remarks
We analyze Heaps’ law through multiple lenses, finding the statistical law’s scaling appears in naturalistic and fictional conversations and that the behavior differs by parts of speech. These differences suggest that statistical laws can inform meaning construction differently through preferencing the use some parts of speech per modality given the modality’s purpose and constraints. Media-specific constraints shape these differences, so it is important to understand how we communicate and connect (engage theory of mind) across technological contexts. The various media pathways for communication are inherently unique in structure and usage, shaping conversation within each modality, and potentially reinforcing thought and behaviors within and between communities. We should ask how those affordances impact our need for social connection [baumeister1995needtobelong, benus2014social_entrainment, sun2020_wellbeing, wheatley2024] and how they support or detract from the social health of our communities [holt2017soc_connection_priority]. Naturalistic conversational data, fictional conversations with their authored intentions, and other online conversational modalities offer ways to analyze these differences. Overall, conversation is rich with incredible context to study how we make sense of each other and coalesce around meaning under different technological constraints on communication.
Acknowledgments
We are grateful for discussion with Alejandro Ruiz, Sarah Phillips, Carter Ward, Fitz Keenan-Koch, Bradford Demarest, and Ben Cooley. Our work was supported in part by MassMutual, National Science Foundation Award #242829 (Science of Online Corpora, Knowledge, and Stories), and a philanthropic gift from an anonymous source.
Appendix A Appendices
Appendix A1 Descriptive Statistics
The normal distribution of conversation scaling provides perhaps a normative finding for naturalistic conversations and could be studied further across naturalistic samples. These distributions are from conducting one regression per conversation assuming the conversation was long enough to be included; thus, Movies (I)’s range is very large. Corpus ranges are: CANDOR: $[0.54,0.81]$, Movies (I): $[0,1.78]$, and Movies (G): $[0.57,0.97]$.
[…]
Figure A1: Distributions of scaling exponents by corpus. Corpus conversation NNs are CANDOR =1,455=1,455, Movies (individual) =3,104=3,104, Movies (grouped) =589=589.
Plots in Figure 2 show the variable nature of each corpus and how we need to carefully treat each as ‘conversation.’ We report additional statistics here per corpus in Table A1 that qualify the spread of the data.
CANDOR Movies (I) Movies (G)
Utterances $36/182/574$ $10/12/89$ $6/473/1286$
Words 2131/5256/162131/5256/16k $21/135/1157$ 61/4833/1461/4833/14k
Speaker words 519/4331/40519/4331/40k $5/108/2587$ $1/135/5911$
Table A1: Raw descriptive statistics of utterances and words by corpus and words by speaker. Values are reported as: ‘min / median / max’.
The next two plots (Figs. A2 and A3) respectively show total and unique conversation count distributions by corpus. Table A2 lists total (T) and unique (U) means and standard deviations, describing the raw data per POS per corpus. Figure 3 normalizes the proportion of each POS to give a better sense of their representation across conversation modes.
CANDOR Movies (I) Movies (G)
Noun (T) 352.24 31.75 1082.45
(141.33) (15.61) (507.24)
(U) 114.70 19.07 292.18
(36.59) (8.24) (108.74)
Verb (T) 326.57 25.70 861.26
(129.92) (13.22) (408.65)
(U) 114.46 19.71 242.45
(29.61) (8.54) (73.45)
Other (T) 622.51 40.77 1399.96
(219.72) (18.84) (650.18)
(U) 82.99 17.98 167.98
(21.15) (6.50) (50.28)
Func (T) 132.51 8.60 276.52
(58.38) (5.84) (139.56)
(U) 22.03 5.85 30.07
(3.94) (2.96) (5.90)
Intj (T) 180.96 3.68 105.29
(61.94) (2.99) (80.23)
(U) 19.38 2.61 21.01
(3.75) (1.61) (10.13)
Table A2: Raw part of speech descriptive statistics by corpus for total words (T) and unique words (U). Values are $\mu$ on top with ($\sigma$) below its mean.
[…]
Figure A2: Part of speech conversation total count distributions by corpus. Each column is respectively CANDOR, Movies (I), and Movies (G).
[…]
Figure A3: Part of speech conversation unique count distributions by corpus. Each column is respectively CANDOR, Movies (I), and Movies (G).
A1.1 Top interjections by corpus
Within corpus, we calculate the most frequent interjections tagged using Stanza. After sorting on CANDOR, we use each corpus’s interjection total to calculate percentages: CANDOR: $299,491$, Movies (I): $9,812$, Movies (G): $61,913$. Any proportions less $0.01$ indicate so while zeros are indicated with a ‘-’. See Table A3 for the top 25 per dataset. The top frequencies demonstrate that the part of speech tagging model and is not without error (tagging some different forms of parallel speech as interjections), but it also shows some obvious differences in type of speech between corpora. Notably, the more space-filling words in English are tagged in naturalistic speech (‘like’, ‘uh’, ‘um’) while we see more tags in fictional speech suggesting the introduction of concepts or contradiction (‘well’); additionally, movies show much higher use of ‘yes’ and ‘no’.
text Candor Movies (I) Movies (G)
yeah 0.35 0.12 0.1
oh 0.11 0.1 0.1
like 0.1 0.01 <0.01<0.01
uh 0.07 0.01 0.02
mhm 0.07 - -
um 0.06 <0.01<0.01 -
ok 0.04 <0.01<0.01 -
no 0.04 0.17 0.14
well 0.03 0.13 0.11
ok. 0.02 - -
wow 0.02 <0.01<0.01 <0.01<0.01
yes 0.02 0.12 0.1
$m$ 0.01 - -
huh <0.01<0.01 0.02 0.02
mhm. <0.01<0.01 - -
hm <0.01<0.01 - -
hello <0.01<0.01 0.01 0.01
yes. <0.01<0.01 - -
hi <0.01<0.01 <0.01<0.01 0.01
$m$. <0.01<0.01 - -
hey <0.01<0.01 0.02 0.03
gosh <0.01<0.01 - -
man <0.01<0.01 <0.01<0.01 <0.01<0.01
no. <0.01<0.01 - -
yep <0.01<0.01 - -
ah - <0.01<0.01 <0.01<0.01
alright - <0.01<0.01 <0.01<0.01
boy - <0.01<0.01 <0.01<0.01
c’mon - <0.01<0.01 <0.01<0.01
damn - - <0.01<0.01
god - - <0.01<0.01
mr - 0.01 0.02
okay - 0.03 0.04
please - 0.02 0.02
shit - <0.01<0.01 <0.01<0.01
sorry - - <0.01<0.01
uhhuh - <0.01<0.01 <0.01<0.01
yah - <0.01<0.01 -
Table A3: Top $25$ interjection frequencies by corpus, calculated as the top appearing within each corpus and divided by its respective total interjection count.
A1.2 CANDOR repeat participants
Last, we include more information regarding repeat speakers in CANDOR. The corpus originally had $1,456$ unique participants, or $1,455$ after we removed one erroneous conversation (technical difficulties). We qualitatively explore the data visually and by conversation length using allotaxonographs [dodds2023allotaxonometry, stonge2025pyallotax]. This exploration suggests some regularities in different topics pertaining to work and locations, which could be an artifact of repeat speakers—we do not conclude that this is problematic for conclusions regarding Heaps’ law but include further quantification of repeat speakers in case it can limit future research. For instance, repeating conversations with strangers could introduce a practice effect for participants choosing to rely on topics and answers they had previously discussed as a method to easily engage in conversation with a stranger. We quantify the number of participants in more than $n$ conversations as $c_{>0}=1,455$: $c_{>1}=754$, $c_{>2}=487$, $c_{>3}=301$, $c_{>4}=185$, $c_{>5}=94$, $c_{>6}=31$, and $c_{>7}=3$.
We refrain from mentioning specific content as researchers must register a request to access teh data. Further investigation of two topics found in the text showed that even of those 2 highly-mentioned topics, the 2 participants making those mentions the most only made them 20 and 26 times in 6 and 4 conversations respectively. The two topics appeared each in less than 1% of utterances in CANDOR. At the maximum, an outsized impact of speakers and a practice effect could look like 3 speakers mentioning the same topics in 8 conversations at a higher rate.
Appendix A2 Variation in Novelty
Complementary measures can show the introduction of information similarly to Heaps’ law. This exploration delves into the interarrival times of unique types and POS, measuring distributions as well as burstiness and memory. Interarrival times reflect the differences within index of occurrences of words and POS categories. For example, in the sequence:
[INTJ, FUNC, INTJ, FUNC, VERB, VERB, NOUN, FUNC, …], starting from the first function word occurrence (FUNC), interarrival times of 2 and 4 reflect differing POS until the next interjection appears (inclusive). Next we plot the interarrival gap distributions by corpus and section of text (Figure A4) as well as by POS (Figure A5).
We offer two methodological notes. First consider documents lengths: Movies (I) distributions are artificially shorter because the text lengths are much shorter. Document length also affects seeing extremely high arrival times of unique types because longer texts have more time for novel types to occur. Second, calculating burstiness requires more than 2 items of the same kind because 2 items provides only 1 interval that results in mean and standard deviation of 0 and a burstiness value of -1, stacking extremely rare words in this region. We filter these rare word occurrences when calculating at the corpora level (on words), but this filtering inherently does not happen for POS because there are 5 categories that are filled every time and do not have rare types. Thus, the POS and corpora-level B and M averages are slightly reflective of this difference while the corpora and literary cluster may be easier to compare.
[…]
Figure A4: Interarrival times by corpus and narrative third in log10 scale.
[…]
Figure A5: Interarrival times by corpus and part of speech in log10 scale. The left column is total tokens while the right column is unique tokens.
CANDOR Movies (I) Movies (G)
Noun .31 .35 .51
Verb .28 .01 .20
Other .24 .60 .14
Func .08 .02 .04
Intj .08 .01 .11
Table A4: Part of speech proportions for the highest interval words by corpus (tail-end of interarrival distributions). The log10 cut off values were respectively [4.2, 2, 4.2] for CANDOR, Movies (I), and Movies (G).
[…]
CANDOR Movies (I) Movies (G)
Noun -.16 / .02 -.29 / -.03 -.26 / .03
Verb -.07 / -.04 -.18 / -.10 -.13 / -.04
Other -.21 / .01 -.26 / -.06 -.24 / -.03
Func .02 / .02 -.23 / -.18 -.03 / .02
Intj .05 / .02 -.43 / -.30 .06 / .03
Figure A6: Burstiness and memory by corpora and parts of speech. Top: Burstiness and memory by corpora and parts of speech. Left is the full metric range from -1 to 1, and the right plot zooms into the data range. Asterisked labels are those corpora shuffled. Bottom: Burstiness and memory by corpora and part of speech with values reported as ‘B / M’. Please refer to the methodological note in this section about comparing corpora averages and POS.
Appendix A3 Methodology Expansion
A3.1 Data cleaning
We recode raw POS tags using the following schema [chacoma_tags_2020]: adverb $\rightarrow$ verb, pronoun $\rightarrow$ noun, proper noun $\rightarrow$ noun, adposition $\rightarrow$ function words, coordinating conjunction $\rightarrow$ function words, subordinating conjunction $\rightarrow$ function words, interjections remained interjections, and all remaining tags $\rightarrow$ other.
A3.2 Regressions
For corpus-level scaling regimes, we use log-space values of $2.0<x\leq 3.4$.
With POS as rows (in order of Noun, Verb, Other, Func, Intj), and corpora as columns (in order of CANDOR, Movies (I), and Movies (G)), log-space start values are:
$\begin{bmatrix}1.4&1.0&1.4\\ 1.5&1.3&1.5\\ 1.5&1.45&1.45\\ 1.25&1.15&1.25\\ 1.2&0.9&1.4\end{bmatrix}$
end values are:
$\begin{bmatrix}3.2&3.2&3.2\\ 3.0&3.0&3.2\\ 3.0&2.6&3.2\\ 1.9&1.9&1.9\\ 2.1&2.2&2.4\end{bmatrix}$
We calculate average exponents over conversations for corpus-level findings. Initially our method involved one regression per conversation, but this approach disqualified numerous very short movie conversations from the corpus-level analysis. As a second approach, we initialize a matrix to be the maximum length of the longest conversation in a corpus, in which we store the number of unique types per conversation. Then, we calculate the average unique types at each step of $N$ up to the maximum based on the number of conversations having types at that ‘window’, excluding zeros from the calculation (where conversations end). This approach works for Movies (I)’s very short conversations, so for comparability of approach, we apply this method across all corpora. This methodological choice makes little difference in the larger corpora’s coefficients:
corpus: $\beta_{\text{former method}}\rightarrow\beta_{\text{final method}}$.
CANDOR: $0.66\rightarrow 0.65$.
Movies (I): $0.76\rightarrow 0.64$.
Movies (G): $0.67\rightarrow 0.63$.
Figure A7 shows conversation-level regressions. We use a stand-out color per corpus to denote outliers where the conversation’s innovation flattened for a minimum raw count of unique types in the conversation ($125$ for CANDOR and Movies (G) and $60$ for Movies (I)). While steps or plateaus appear in the lower-left quadrant on these charts, these visual steps do not meet the threshold for outliers. These areas visually look like steps because raw data transformed to $log_{10}$ scale produce tiny decimal increments.
[…]
Figure A7: Conversation-level regressions by corpora. These plots show one regression per conversation, an alternative method we decided against in the current project. In A) CANDOR, B) Movies (I), and C) Movies (G), the stand-out line is an outlier conversation.
Appendix A4 CANDOR’s Naturalistic Features
In this work, we could not compare CANDOR’s rich naturalistic features to other corpora as these features can not be engineered from transcripts alone. Through CANDOR only, we explore some of these features.
A4.1 Pauses
We find some erroneously recorded start/stop times that result in implausible pauses (approximately 25 minutes). These could be identified because the recording start/time time jumps for just that utterance or the pause does not reflect the start/stop time difference. At least 6 conversations were affected at a minimum.
We remove these few data from analysis. Second, in qualitatively exploring longer-than-expected pause ranges (e.g., over 200 seconds), we find evidence of numerous video chat connection issues in participants’ conversations, but it remains unclear at what range we start to see substantive connection issues. We are able to easily identify and remove 8 erroneous pauses, but future work should systematically classify and restrict these occurrences from analysis. A method could simply be outlier identification and removal bolstered by language suggestive of technology problems.
Note that the range for pauses (in seconds) is negative to positive, reflecting overlap in speech (interruptions) or otherwise gaps between speakers’ turns. Some utterances have a list of pauses which seems to reflect the timings around a listener’s parallel speech. These are not representative of pauses’ overlaps (interruptions) or gaps between turns, so we remove these pauses for reporting. In fact, with these pauses included, the average is negative ($\mu=-0.22,~\text{median}=1.18,~\sigma=15.24$) but positive with parallel speech pauses removed ($\mu=1.14,~\text{median}=0.61,~\sigma=8.04$). We find that these statistics are sensitive to the technical problems’ outliers, hence including the median. Prior CANDOR work [reece2023candor] using the Backbiter model showed the average pause interval was 0.2 seconds ($\text{median}=0.38$ seconds) with gaps largely under 1 second. The difference between NaturalTurn and Backbiter is that NaturalTurn keeps a speaker’s turn together while preserving concurrent parallel speech, but the parallel speech ‘interruptions’ do not intersperse the speaker’s coherent turn; thus, the length of utterances and pauses may differ from those reported under Backbiter. Given that online connection issues seem present in the turn segmenting, it tended to stretch the mean and standard deviation, so we suggest more reliance on the median and methods robust to outliers.
A4.2 Conversation-level features
Next are conversation-level distributions for backchannels, secondary speech, secondary + backchannel occurrences, and pauses. We normalize these occurrences by number of utterances in the conversation. We quantify the spread of spread of these parallel speech features, which could inform future work related to parallel speech; parallel speech includes language that mirror’s a storyteller’s emotion or reinforces key moments [cooney2024naturalturn]—both of which can be pro-social in nature.
We make total utterances the unit of conversation length because parallel speech typically seems to require a coherent thought/idea for there to be parallel speech expressed. Thus, the pauses’ unit range up to $2.0$ makes more sense than when normalized by words—much of which may not merit a listener’s parallel speech. We create a research-informed schema of scoring the expected number of pauses (gap time between speaker turns). Other work [templeton2023long_gaps] defined long gaps as $2$ seconds which signaled awkwardness in the conversation.
Using these descriptions, we score pauses using this respective schema: 1,0,1,2-1,0,1,2 if pauses are less than 0, less than $0.4$ seconds, less than $4$ seconds, and equal to or above $1.25$ seconds. These scores are an exploration into what normative pauses and gaps may feel like in conversation.
When pause scores in a conversation are normalized by the amount of utterances, the distribution becomes normal (counter to its skew in both raw amount of pauses and unnormalized pause scores per conversation). This normal appearance may support some adherence to pauses that feel more tenable or desirable in a conversation. Further work should connect scores such as these to quality outcomes. It would also be interesting to asses response gaps in non-naturalistic conversation mediums such as online to see if the same perception exists regarding time-to-reply gaps.
Backchannels, secondary speech, and pauses are rich in interpersonal connection information, especially regarding interjectional phrases. This data may be useful when trying to establish a baseline of interjections that are not only form-based but context-based as well considering that Stanza seems unable to find these naturalistic features in the way that the NaturalTurn model can.
See Figure A8 for parallel speech and pause distributions when normalized at the conversation level. A conversation’s value was divided by its count of utterances.
[…]
$\mu$ $\sigma$ median
Backchannels .38 .22 .33
Secondary .11 .07 .10
All parallel .50 .26 .45
Pause score .80 .17 .79
Figure A8: Top: Parallel (left column) and pause (right column) variables at the conversation level, normalized by a conversation’s count of utterances. Bottom: Descriptive statistics for CANDOR’s parallel speech and pause variables at the conversation level, normalized by length of conversations (utterances).